# Model-Independent Machine Learning Approach for Nanometric Axial Localization and Tracking

### v1

@@ Abstract
Accurately tracking particles and determining their position along the optical axis is a major challenge in optical microscopy, especially when extremely high precision is needed. In this study, we introduce a deep learning approach using convolutional neural networks (CNNs) that can determine axial positions from dual-focal plane images without relying on predefined models. Our method achieves an axial localization accuracy of 40 nanometers—six times better than traditional single-focal plane techniques. The model’s simple design and strong performance make it suitable for a wide range of uses, including dark matter detection, proton therapy for cancer, and radiation protection in space. It also shows promise in fields like biological imaging, materials science, and environmental monitoring. This work highlights how machine learning can turn complex image data into reliable, precise information, offering a flexible and powerful tool for many scientific applications.
@@ 1 Introduction
One such application is directional dark matter search experiments that require precise measurements of ions recoiling after their interactions with dark matter particles [1, 2]. Due to their extremely low kinetic energies, in the $1\div 100$ keV range, recoiling ions produce tracks ranging from a few millimeters in gases at low pressure to a few hundreds of nanometers in solids [2, 3]. Taking into account that the required detector mass in practice amounts to several tons, the choice of solid materials as a sensitive medium is unavoidable.
The only detector technology that has shown so far to be capable of reconstructing sub-micrometer recoil tracks in a solid medium is the nuclear emulsion [4]. It has been demonstrated that it is possible to reconstruct tracks with lengths down to 50 nm using special emulsion films with nanometric crystals and a custom-designed super-resolution optical microscope [5]. However, the super-resolution is achievable only in the transverse plane while in the axial direction the resolution is still diffraction-limited. Enabling the super-resolution also in the axial direction will make it possible to achieve a precise 3D measurement of recoiling ion tracks, thus improving the overall detector sensitivity to dark matter by almost one order of magnitude [6].
@@
One possible way to overcome the diffraction-limited axial localization and axial resolution problems is to use the multifocal plane imaging technique that simultaneously captures two separate focal planes within the specimen, providing extra information [12]. However, this method relies on complex mathematical models to describe light interactions and propagation, which often involve numerous parameters that are difficult to determine precisely.
This study introduces a model-independent, deep learning-based approach for nanometric axial localization and particle tracking using convolutional neural networks (CNNs) applied to dual-focal plane images. Traditional techniques often rely on detailed system parameters and are hard to adapt to different imaging setups, limiting their use in real-time or cross-disciplinary applications. The novelty lies in that our model learns the relationship between image features and axial coordinates directly from image data. This eliminates the need for detailed knowledge of the optical system or the physical properties of the objects under study. Moreover, the model-independent nature of this technique makes it highly adaptable, allowing it to be applied to various imaging systems without requiring significant adjustments. This feature makes the method particularly valuable for interdisciplinary applications where the underlying imaging conditions or object properties may vary significantly.
@@ 2.1 Image dataset
The images for this study were taken with a high-resolution, single focal plane optical microscope described in the Appendix section. The scanned sample contained silver nanoparticles with a diameter of 60 nm randomly immersed in gelatin. The whole thickness of 20 $\mu$m was sampled with steps of 250 nm, a value chosen independently of this study, based on prior research demonstrating that it provides a balance between axial resolution and data acquisition speed. Images of nanoparticles could appear anywhere in the field of view and were not necessarily in focus. Then the nanoparticle images, both focused and unfocused, were detected by a program, cropped, and saved to files along with the corresponding X, Y, and Z coordinates.
Figure 1 shows an example of images of a single nanoparticle taken at different defocus values $\Delta{z}$. As a result of minor uncorrected spherical aberration, the Airy pattern is clearly visible predominantly in the negative $Z$ direction. In contrast, in the positive $Z$ direction, approaching the objective lens, the rings appear blurred. This asymmetry also causes the shape of the brightness profile of the nanoparticles, depicted in Figure 2, to differ slightly from a Gaussian distribution shown as a red dotted line. However, our deep learning approach is data-driven and does not assume a specific model, allowing it to adapt to these variations and maintain high localization accuracy.
Figure 1: Images of the same nanoparticle acquired at different depths with step of $\sim$250 nm. The dimensions of the image are $2150\times 2150$ nm. The $Z$ axis is directed towards the objective lens.
Figure 2: Mean brightness profile along the optical axis averaged over all nanoparticles. The $Z$ axis is directed towards the objective lens.
@@ 2.2 Nanoparticle reference axial coordinate definition
The vertical sampling step of the image dataset used for this study is approximately $250$ nm, ensuring that a nanoparticle appears in at least three consecutive images. This allowed for the calculation of the axial position using the weighted average of the brightness distribution:
@@
The axial localization accuracy is defined by:
@@
where $\sigma_{z}^{im}=35$  nm is the vertical stage linear encoder accuracy and $\sigma_{B}^{im}=0.73$ is the noise at the videocamera sensor, calculated as the RMS pixel brightness variation. The mean localization accuracy determined by the brightness profile using equation 2.2 is $\sigma_{z}^{prof}\equiv\bar{\sigma}_{\bar{z_{i}}}=(18\pm 1)$ nm averaged over all nanoparticles.
The value $\bar{z_{i}}$ calculated using equation 2.1 is used in training as the reference axial coordinate of nanopaticle ${i}$. Each calculation involves several (typically 5-6) consecutive images taken at different depths by a single focal plane microscope. While this approach is effective for measuring static objects, it is not practical for real-time applications. Capturing multiple images sequentially requires time for stage movement and stabilization, making it too slow for dynamic objects. Consequently, a single focal plane microscope cannot benefit from multiple-depth image fitting, and its axial localization accuracy remains limited by the width of the brightness profile shown in Figure 2.
@@ 2.3 Training datasets
This study aims to develop a novel measurement approach tailored for multifocal plane microscopes. To replicate multiple focal planes, we selected images from the existing dataset to form image pairs that are offset by $\delta$ relative to each other. For each pair of images, the defocus distance from the corresponding nanoparticle to the bottom image (with the smaller $Z$ coordinate) in the pair is calculated as $\Delta{z}_{k}=z_{k}^{bottom}-\bar{z}_{i}$, where $z_{k}^{bottom}$ is the axial coordinate of the bottom image in the pair $k$ and $\bar{z}_{i}$ is the axial coordinate of the nanoparticle $i$ to which the image pair $k$ belongs. By applying this procedure to all nanoparticles, we establish a set of image pairs $k$ along with the corresponding defocus distance $\Delta{z}_{k}$. Such a triplet: two images offset by $\delta$ and the defocus distance $\Delta{z}_{k}$, form an independent measurement carried out by a dual-focal plane microscope.
We denote as ${\Delta{z}_{k}}^{test}$ the defocus distances calculated using $\bar{z}_{i}$ measured with the brightness profile and as ${\Delta{z}_{k}}^{pred}$ those predicted by CNN. Values of ${\Delta{z}_{k}}^{test}$ are used for training. During the test phase, CNN predicts ${\Delta{z}_{k}}^{pred}$ that are later compared with $\Delta{z}_{k}^{test}$ to estimate the localization accuracy.
@@ 2.4 CNN model architecture and training
As reported in Figure 3, the implemented CNN for our study is composed of three convolutional layers with size 16, 32 and 64, respectively. The convolution operation is performed with a filter of $3\times 3$ by using as padding the same strategy. The output of the convolution operation is passed through a nonlinear activation function. In this study, the most common nonlinear activation function, namely Rectified Linear Unit (ReLU), is used. This function returns either the same value given in input, or zero if the input value is negative. Formally,
@@
The output of the activation function is undergone to a Batch Normalization (BN) layer that applies a transformation that maintains the mean output close to 0 and the output standard deviation close to 1. Finally, the pooling layer downsamples the input representation by taking, in our case, the average value over the window defined by pool size ($2\times 2$) for each dimension along the features axis. In order to given in input data to the two fully connected layers, a flatten layer is necessary to rearrange the three dimensional data, obtained in output from the last pooling layer, in a one dimensional vector. In particular, the last fully connected layer is added to perform regression. Therefore, the output of the first fully connected layer is passed through a ReLU activation function and a BN layer, whereas, the output of the second fully connected layer is passed through a linear activation function to perform regression. Moreover, a dropout layer is added in order to randomly set input units to 0 (removing neurons) with a frequency of rate that is equal to 0.5 between the two fully connected layer in order to prevent overfitting.
@@ 3.1 Deep Learning approach
To achieve high-accuracy axial localization, we implemented a CNN-based regression model designed to predict axial coordinates from dual-focal plane image pairs. The network architecture consists of:
@@
The CNN model transforms raw image pairs into axial coordinate predictions by learning implicit mappings between pixel intensity patterns and object depth. The extracted axial coordinates represent actionable knowledge, enabling precise 3D localization without explicit physical models. This data-driven approach offers a generalizable solution adaptable to diverse imaging systems and scientific applications.
@@
Offset $\delta$, nm Image pairs Localization accuracy, nm
250 33562 $48.4\pm 0.7$
500 34842 $40.0\pm 0.5$
750 21685 $44.0\pm 0.7$
Figure 4: (a) Predicted value ${\Delta{z}_{k}}^{pred}$ as a function of test value ${\Delta{z}_{k}}^{test}$ for image offset $\delta=500$ nm. Prediction accuracy $({\Delta{z}_{k}}^{pred}-{\Delta{z}_{k}}^{test})$ for image offsets $\delta$ equal to (b) 500 nm, (c) 250 nm and (d) 750 nm.
@@ 3.2 Measurement results and performance
After training, CNN performance was tested on a fraction of the image dataset that was not used for training ($\sim{20}\%$ of the entire dataset). The plot in Figure LABEL:fig:3_d500 shows the proportionality of the predicted defocus values to the expected ones, as well as the independence of performance from the defocus value for image offset $\delta=500$ nm. For image offsets $\delta=250$ nm and $\delta=750$ nm the corresponding plots are similar. The periodic structure in the proportionality plot that repeats every $250$ nm is due to the image acquisition scheme in which the images are taken with the same sampling step of $250$ nm.
The difference between the predicted defocus values ${\Delta{z}_{k}}^{pred}$ and the calculated ones ${\Delta{z}_{k}}^{test}$ defines the precision of the CNN and is shown in Figure LABEL:fig:4_d500,  LABEL:fig:4_d250 and Figure LABEL:fig:4_d750 for image offsets $\delta=500$ nm, $\delta=250$ nm and $\delta=750$ nm, respectively. The axial localization accuracy $\sigma_{z}$ is defined by the width of the difference distribution $({\Delta{z}_{k}}^{pred}-{\Delta{z}_{k}}^{test})$ and is summarized in the third column of Table 1 for each image dataset.
The red dotted line in Figure 2 represents the Gaussian fitting of the brightness profile. The width of the fit allows for a rough estimate of the axial localization accuracy of the single focal plane microscope used for image acquisition to be approximately $\sigma_{z}^{single}\approx{260}$ nm.
In case when two images are used in combination with the reported approach, the axial localization accuracy improves significantly and reaches $\sigma_{z}^{\delta=500}=40$ nm for the offset $\delta=500$ nm. Somewhat lower performance for offsets $\delta=250$ nm and $\delta=750$ nm implies that offset $\delta=500$ nm is closer to the optimal value, whose existence is demonstrated in ref. [16]. To put it simply, images with a distance of $\delta=250$ nm are too close and their differences are insufficient to provide the best localization accuracy. In contrast, images with a distance of $\delta=750$ nm are too far apart. In this case, the brightness of one image is often too weak, leading to the loss of details due to a too low signal-to-noise ratio.
@@ 4 Discussion
The proposed approach enables the instantaneous measurement of the axial coordinate of a nanoparticle without requiring refocusing or capturing a stack of images at different depths. The microscope setup requires at least two focal planes offset by a few hundred nanometers. This approach can be easily extended to calculate all three spatial coordinates or to incorporate more than two focal planes, enhancing its flexibility. Moreover, it does not involve explicit fitting to point spread function (PSF) models, making it applicable to complex objects with nontrivial optical properties.
For this study, we utilized an image dataset obtained with a single focal plane microscope by sequential in-depth image acquisition at a step of approximately 250 nm. To simulate a multifocal plane microscope, we selected image pairs spaced by specific distances: 250, 500, or 750 nm. The nanoparticles were immobilized in gelatin, ensuring negligible displacement during image acquisition. However, mechanical motion during image capture introduced small variations in focal plane distance, which impacted the precision of CNN predictions. Despite this, the achieved axial localization accuracy of 40 nm, as shown in Figure LABEL:fig:4_d500, represents a six-fold improvement over traditional single focal plane microscopes (see Figure 2).
@@
This study highlights the capability of deep learning to surpass conventional limitations in axial localization, achieving nanometric precision with broad interdisciplinary implications. While the immediate applications include dark matter detection, proton therapy, and space radiation protection, the methodology is highly generalizable. By providing a robust, transparent, and scalable framework, this approach paves the way for future innovations in super-resolution imaging, knowledge extraction, and real-time nanometric tracking across multiple scientific and technological domains.

### v2

@@ Abstract
Accurately tracking particles and determining their coordinate along the optical axis is a major challenge in optical microscopy, especially when extremely high precision is needed. In this study, we introduce a deep learning approach using convolutional neural networks (CNNs) that can determine axial coordinates from dual-focal-plane images without relying on predefined models. Our method achieves an axial localization precision of 40 nanometers—six times better than traditional single-focal-plane techniques. The model’s simple design and strong performance make it suitable for a wide range of uses, including dark matter detection, proton therapy for cancer, and radiation protection in space. It also shows promise in fields like biological imaging, materials science, and environmental monitoring. This work highlights how machine learning can turn complex image data into reliable, precise information, offering a flexible and powerful tool for many scientific applications.
@@ 1 Introduction
One such application is directional dark matter search experiments that require precise measurements of ions recoiling after their interactions with dark matter particles [1, 2]. Due to their extremely low kinetic energies, in the 11001-1001 - 100 keV range, recoiling ions produce tracks ranging from a few millimeters in gases at low pressure to a few hundreds of nanometers in solids [2, 3]. Taking into account that the required detector mass in practice amounts to several tons, the choice of solid materials as a sensitive medium is advantageous.
The only detector technology that has shown so far to be capable of reconstructing sub-micrometer recoil tracks in a solid medium is the nuclear emulsion [4]. It has been demonstrated that it is possible to reconstruct tracks with lengths down to 50 nm using special emulsion films with nanometric crystals and a custom-designed super-resolution optical microscope [5]. In general, super-resolution refers to any imaging technique (optical, computational, or hybrid) that achieves spatial resolution beyond the classical diffraction limit of light, typically around 200–250 nm in the transverse plane and 500–700 nm in the axial direction for visible wavelengths. In the context of the study reported in [5], super-resolution is achieved in the transverse plane, where the system reaches resolutions of approximately 50–100 nm. However, the axial resolution remains diffraction-limited. Enabling the super-resolution also in the axial direction will make it possible to achieve a precise 3D measurement of recoiling ion tracks, thus improving the overall detector sensitivity to dark matter by almost one order of magnitude [6].
@@
One possible way to overcome the diffraction-limited axial localization and axial resolution problems is to use the dual-focal-plane imaging technique that simultaneously captures two separate focal planes within the specimen, providing extra information [12]. However, this method relies on complex mathematical models to describe light interactions and propagation, which often involve numerous parameters that are difficult to determine precisely.
This study introduces a model-independent, deep learning-based approach for nanometric axial localization and particle tracking using convolutional neural networks (CNNs) applied to dual-focal-plane images. Traditional techniques often rely on detailed system parameters and are hard to adapt to different imaging setups, limiting their use in real-time or cross-disciplinary applications. The novelty lies in that our model learns the relationship between image features and axial coordinates directly from image data. This eliminates the need for detailed knowledge of the optical system or the physical properties of the objects under study. Moreover, the model-independent nature of this technique makes it highly adaptable, allowing it to be applied to various imaging systems without requiring significant adjustments. This feature makes the method particularly valuable for multidisciplinary applications where the underlying imaging conditions or object properties may vary significantly.
@@ 2.1 Image dataset
The images for this study were taken with a high-resolution, single-focal-plane optical microscope described in the Appendix section. The scanned sample contained silver nanoparticles with a diameter of 60 nm randomly immersed in gelatin. The whole thickness of 20 $\mu$m was sampled with steps in the range 245-255 nm, a value chosen independently of this study, based on prior research [13] demonstrating that it provides a balance between axial resolution and data acquisition speed. Images of nanoparticles could appear anywhere in the field of view and were not necessarily in focus. Then the nanoparticle images, both focused and unfocused, were detected by a program, cropped, and saved to files along with the corresponding X, Y, and Z coordinates.
Figure 1 shows an example of images of a single nanoparticle taken at different defocus values $\Delta{z}$. As a result of minor uncorrected spherical aberration, the Airy pattern is clearly visible predominantly in the negative $Z$ direction. In contrast, in the positive $Z$ direction, approaching the objective lens, the rings appear blurred. This asymmetry also causes the shape of the brightness profile of the nanoparticles, depicted in Figure 2, to differ slightly from a Gaussian distribution shown as a red dotted line. However, our deep learning approach is data-driven and does not assume a specific model, allowing it to adapt to these variations and maintain high localization precision.
Figure 1: Images of the same nanoparticle acquired at different axial coordinates with step in the range 245255245-255245 - 255 nm. The dimensions of the image are $\times 2150$ 2150 nm. The $Z$ axis is directed towards the objective lens.
Figure 2: Mean brightness profile along the optical axis averaged over 105913 nanoparticles. Error bar values are calculated as the RMS of the nanoparticle brightness distribution in the corresponding bin. The $Z$ axis is directed towards the objective lens.
@@ 2.2 Nanoparticle reference axial coordinate definition
The sampling step in the axial direction of the image dataset used for this study is in the range 245255245-255245 - 255 nm, ensuring that a nanoparticle appears in at least three consecutive images. This allowed for the calculation of the axial coordinate using the weighted average of the brightness distribution:
@@
The axial localization precision is defined by:
@@
where $\sigma_{z}^{im}=35$  nm is the objective lens coordinate measurement precision and $\sigma_{B}^{im}=0.73$ is the noise at the videocamera sensor, calculated as the RMS pixel brightness variation. The mean localization precision determined by the brightness profile using equation 2.2 is $\sigma_{z}^{prof}\equiv\bar{\sigma}_{\bar{z_{i}}}=(18\pm 1)$ nm averaged over 105913 nanoparticles.
The value $\bar{z_{i}}$ calculated using equation 2.1 is used in training as the reference axial coordinate of nanopaticle ${i}$. Each calculation involves several (typically 5-6) consecutive images taken at different axial coordinates by a single-focal-plane microscope. While this approach is effective for measuring static objects, it is not practical for real-time applications. Capturing multiple images sequentially requires time for stage movement and stabilization, making it too slow for dynamic objects. Consequently, a single-focal-plane microscope cannot benefit from dual-focal-plane image fitting, and its axial localization precision remains limited by the width of the brightness profile shown in Figure 2.
@@ 2.3 Training datasets
This study aims to develop a novel measurement approach tailored for dual-focal-plane microscopes. To replicate two focal planes, we selected images from the existing dataset to form image pairs that are offset by $\delta$ relative to each other. For each pair of images, the defocus distance from the corresponding nanoparticle to the bottom image (with the smaller $Z$ coordinate) in the pair is calculated as $\Delta{z}_{k}=z_{k}^{bottom}-\bar{z}_{i}$, where $z_{k}^{bottom}$ is the axial coordinate of the bottom image in the pair $k$ and $\bar{z}_{i}$ is the axial coordinate of the nanoparticle $i$ to which the image pair $k$ belongs. By applying this procedure to all nanoparticles, we establish a set of image pairs $k$ along with the corresponding defocus distance $\Delta{z}_{k}$. Such a triplet: two images offset by $\delta$ and the defocus distance $\Delta{z}_{k}$, form an independent measurement carried out by a dual-focal-plane microscope.
We denote as ${\Delta{z}_{k}}^{test}$ the defocus distances calculated using $\bar{z}_{i}$ measured with the brightness profile and as ${\Delta{z}_{k}}^{pred}$ those predicted by CNN. Values of ${\Delta{z}_{k}}^{test}$ are used for training. During the test phase, CNN predicts ${\Delta{z}_{k}}^{pred}$ that are later compared with $\Delta{z}_{k}^{test}$ to estimate the localization precision.
@@ 2.4 CNN model architecture and training
As reported in Figure 3, the implemented CNN for our study is composed of three convolutional layers with size 16, 32 and 64, respectively. The convolution operation is performed with a filter of $\times 3$ 3 by using as padding the same strategy. The output of the convolution operation is passed through a nonlinear activation function. In this study, the most common nonlinear activation function, namely Rectified Linear Unit (ReLU), is used. This function returns either the same value given in input, or zero if the input value is negative. Formally,
@@
The output of the activation function is undergone to a Batch Normalization (BN) layer that applies a transformation that maintains the mean output close to 0 and the output standard deviation close to 1. Finally, the pooling layer downsamples the input representation by taking, in our case, the average value over the window defined by pool size ($\times 2$ 2) for each dimension along the features axis. In order to provide input data to the two fully connected layers, a flatten layer is necessary to rearrange the three dimensional data, obtained in output from the last pooling layer, in a one dimensional vector. In particular, the last fully connected layer is added to perform regression. Therefore, the output of the first fully connected layer is passed through a ReLU activation function and a BN layer, whereas, the output of the second fully connected layer is passed through a linear activation function to perform regression. Moreover, a dropout layer is added in order to randomly set input units to 0 (removing neurons) with a frequency of rate that is equal to 0.5 between the two fully connected layer in order to prevent overfitting.
@@
The training parameters were selected based on a combination of standard practices in deep learning and a dedicated hyperparameter optimization study. A systematic scan of hyperparameters was performed, including batch size (4 to 64), dropout rate (0.4 to 0.6), and number of epochs (30 to 600), while monitoring the resulting localization precision. The batch size of 16 consistently yielded the best trade-off between convergence stability and generalization, while 400 epochs provided sufficient training without overfitting. The learning rate of 0.001 was selected based on standard practice for the Adam optimizer and was found to be effective without requiring further tuning.
@@ 3.1 Deep Learning approach
To achieve high-precision axial localization, we implemented a CNN-based regression model designed to predict axial coordinates from dual-focal-plane image pairs. The network architecture consists of:
@@
The CNN model transforms raw image pairs into axial coordinate predictions by learning implicit mappings between pixel intensity patterns and object axial coordinate. The extracted axial coordinates represent actionable knowledge, enabling precise 3D localization without explicit physical models. This data-driven approach offers a generalizable solution adaptable to diverse imaging systems and scientific applications.
@@
Offset $\delta$, nm Image pairs Localization precision, nm
Test Training
250 33562 48.4±0.748.4$\pm$ 0.748.4 ± 0.7 46.2±0.646.2$\pm$ 0.646.2 ± 0.6
500 34842 40.0±0.540.0$\pm$ 0.540.0 ± 0.5 38.9±0.538.9$\pm$ 0.538.9 ± 0.5
750 21685 44.0±0.744.0$\pm$ 0.744.0 ± 0.7 42.6±0.642.6$\pm$ 0.642.6 ± 0.6
Figure 4: (a) Predicted value ${\Delta{z}_{k}}^{pred}$ as a function of test value ${\Delta{z}_{k}}^{test}$ for image offset $\delta=500$ nm. Prediction precision $({\Delta{z}_{k}}^{pred}-{\Delta{z}_{k}}^{test})$ for image offsets $\delta$ equal to (b) 500 nm, (c) 250 nm and (d) 750 nm.
@@ 3.2 Measurement results and performance
[…]
The difference between the predicted defocus values ${\Delta{z}_{k}}^{pred}$ and the calculated ones ${\Delta{z}_{k}}^{test}$ defines the precision of the CNN and is shown in Figure LABEL:fig:4_d500,  LABEL:fig:4_d250 and Figure LABEL:fig:4_d750 for image offsets $\delta=500$ nm, $\delta=250$ nm and $\delta=750$ nm, respectively. The axial localization precision $\sigma_{z}$ is defined by the width of the difference distribution $({\Delta{z}_{k}}^{pred}-{\Delta{z}_{k}}^{test})$ and is summarized in the third column of Table 1 for each image dataset. These values, obtained from the test set, are compatible with the localization precision achieved on the training data, shown in the fourth column of Table 1. The training precision values are slightly better than the test ones, which is expected due to the reduced variability in the training dataset and the optimization bias introduced during model fitting. In ideal conditions, where the model generalizes well and overfitting is absent, the residual distributions for training and test sets are expected to be similar. However, the test set may contain more diverse examples and is not directly optimized during training, which can lead to a slightly broader residual distribution and thus a marginally higher localization uncertainty.
The red dotted line in Figure 2 represents the fitting of the brightness profile. To model the brightness distribution along the optical axis, we used a Gaussian function to represent the nanoparticle signal and added a constant term to account for the background intensity. This approach, often referred to as a Gaussian fit with offset, allows for more accurate estimation of the peak position by separating signal from baseline noise. The width of the fit allows for a rough estimate of the axial localization precision of the single-focal-plane microscope used for image acquisition to be approximately $\sigma_{z}^{single}\approx{260}$ nm.
In case when two images are used in combination with the reported approach, the axial localization precision improves significantly and reaches $\sigma_{z}^{\delta=500}=40$ nm for the offset $\delta=500$ nm. Somewhat lower performance for offsets $\delta=250$ nm and $\delta=750$ nm implies that offset $\delta=500$ nm is closer to the optimal value, whose existence is demonstrated in ref. [17]. To put it simply, images with a distance of $\delta=250$ nm are too close and their differences are insufficient to provide the best localization precision. In contrast, images with a distance of $\delta=750$ nm are too far apart. In the latter case, the brightness of one of the images is often too weak, leading to the loss of details due to a too low signal-to-noise ratio.
@@ 4 Discussion
The proposed approach enables the instantaneous measurement of the axial coordinate of a nanoparticle without requiring refocusing or capturing a stack of images at different axial coordinates. The microscope setup requires at least two focal planes offset by a few hundred nanometers. This approach can be easily extended to calculate all three spatial coordinates or to incorporate more than two focal planes, enhancing its flexibility. Moreover, it does not involve explicit fitting to point spread function (PSF) models, making it applicable to complex objects with nontrivial optical properties.
For this study, we utilized an image dataset obtained with a single-focal-plane microscope by sequential image acquisition at the axial step in the range 245255245-255245 - 255 nm. To simulate a dual-focal-plane microscope, we selected image pairs spaced by specific distances: 250, 500, or 750 nm. The nanoparticles were immobilized in gelatin, ensuring negligible displacement during image acquisition. However, mechanical motion during image capture introduced small variations in focal plane distance, which impacted the precision of CNN predictions. Despite this, the achieved axial localization precision of 40 nm, as shown in Figure LABEL:fig:4_d500, represents a six-fold improvement over traditional single-focal-plane microscopes (see Figure 2).
@@
This study highlights the capability of deep learning to surpass conventional limitations in axial localization, achieving nanometric precision with broad multidisciplinary implications. While the immediate applications include dark matter detection, proton therapy, and space radiation protection, the methodology is highly generalizable. By providing a robust, transparent, and scalable framework, this approach paves the way for future innovations in super-resolution imaging, knowledge extraction, and real-time nanometric tracking across multiple scientific and technological domains.